The cost of laminate flooring is $5.89 per sq. ft. if Nancy installs it herself or $8.89 per sq. ft. if installed by flooring company. Determine the cost for the flooring in the living room/dining room.
Living/Dining Room is 25' x 22'

Answer by nerdybill(7384)

The cost of laminate flooring is $5.89 per sq. ft. if Nancy installs it herself or $8.89 per sq. ft. if installed by flooring company. Determine the cost for the flooring in the living room/dining room.
Living/Dining Room is 25' x 22'
.
Area of room = 25*22 = 550 square feet
.
Cost if installed by Nancy:
550*5.89 = $3239.50
.
Cost if installed by flooring company:
550*8.89 = $4889.50